OpenStudy (anonymous):

if you borrow $5 today the at the end of the month you will have to pay $6 with this rate of interest COMPOUNDING MONTHLY. supposing that you borrowed $1000 today, how much will be your dues after 6 months if you are unable to pay any amount during such period.

ganeshie8 (ganeshie8):

so you have borrowed `$1000` today, can you tell how much you need to pay after `1` month ?

OpenStudy (anonymous):

$200/month

OpenStudy (anonymous):

based on my computation the interest is 20% per month.

ganeshie8 (ganeshie8):

Exactly !

ganeshie8 (ganeshie8):

may i know how u got $200 ?

OpenStudy (anonymous):

because I borrow $1000 and the interest per month is 20% what i did is this. $1000 X 20% = 200

OpenStudy (anonymous):

is that right?

ganeshie8 (ganeshie8):

thats absolutely right ! another way to work it without using percentages is : for every $5, you're payign $1 interest so for $1000, you would be paying $1000/5 interest

ganeshie8 (ganeshie8):

today : $1000 after `1` month : $1000 + $1000/5 = $1000(1 + 1/5) = $1000(1+ 0.2)

OpenStudy (anonymous):

i see..

ganeshie8 (ganeshie8):

after `2` months : = `$1000(1+ 0.2)` + `$1000(1+ 0.2)`/5 = `$1000(1+ 0.2)`(1+ 1/5) = `$1000(1+ 0.2)`^2

ganeshie8 (ganeshie8):

the pattern continues...

ganeshie8 (ganeshie8):

after `6` months : `$1000(1+ 0.2)`^6

ganeshie8 (ganeshie8):

thats just a compound interest formula ^^

ganeshie8 (ganeshie8):

use ur calculator to evaluate that expression : http://www.wolframalpha.com/input/?i=1000%281%2B0.2%29%5E6

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

the total is 2,985.98

OpenStudy (anonymous):

i thought the formula for compound interest A=P(1+r/n)nt

ganeshie8 (ganeshie8):

yes, thats the formula : P = 1000 r = 0.2 n = 1 t = 6

OpenStudy (anonymous):

may I ask why "n" is 1?

OpenStudy (anonymous):

ok I get it..thank u for answering my question. :)

ganeshie8 (ganeshie8):

np :) n=1 because your interest rate is 20% per MONTH, since we're compounding MONTHly, number of compounding intervals (n) would be just `1`

OpenStudy (anonymous):

yup. that's what you did a while ago. you divide 1/5 and you get .2

OpenStudy (anonymous):

yours is much simple than to my teacher. Thank you so much! and it's very clear to me now!

ganeshie8 (ganeshie8):

glad to hear that :) yw :)

OpenStudy (jmark):

$6-$5 =$1 $1 is 20% of $5=$1 so, 20% of $1000=$200 after 6 months=$1200 CBSE Class 7 Maths Question Papers - http://edworld.hpage.co.in/cbse-class-7-maths-question-papers_3668604.html